# MorphAny3D: Unleashing the Power of Structured Latent in 3D Morphing

### v2

@@ Abstract
Figure 1: MorphAny3D enables smooth, semantically coherent, and aesthetically pleasing transitions from source objects (first column) to target objects (last column), even when they share no semantic or visual relationship (e.g., bee to biplane, Wukong to tree).
@@ 1 Introduction
  • We propose MorphAny3D, the training-free 3D morphing framework based on Structured Latent (SLAT), capable of generating smooth and semantically coherent deformations across diverse 3D object categories.
@@ 3.3 SLAT Fusion Patterns in Attention Mechanism
From these results, we draw three key observations: (1) KV-Fused CA significantly enhances the structural and semantic plausibility of 3D morphing by fusing 2D semantic information from the source and target objects within cross-attention (the lowest FID). However, it fails to eliminate localized implausible artifacts, as shown in the orange box of Fig. 4-(b). (2) KV-Fused SA effectively improves the smoothness and continuity of the morphing sequence by aggregating 3D latent features from the source and target objects within self-attention (the lowest PPL). (3) Yet, simultaneously applying KV-fused CA and KV-fused SA—while aiming to maximize both plausibility and smoothness—compromises the plausibility of the resulting deformation (see Fig. 4-(d)). These results indicate that while SLAT feature blending benefits morphing, naive combination harms the plausibility-smoothness trade-off. To address this, we propose two tailored modules—Morphing Cross-Attention and Temporal-Fused Self-Attention—that preserve the advantages of fusion while avoiding its pitfalls.
Figure 5: Attention maps visualization for different attention mechanisms. Red stars denote SLAT head features; pink stars mark their corresponding input regions. Orange boxes highlight KV-Fused CA’s incorrect attention focus. MCA preserves correct, semantically consistent attention and avoids KV-Fused CA’s artifacts shown in Fig. 4-(b).
@@ 3.4 Morphing Cross-Attention (MCA)
We hypothesize that the local irrational structure in KV-Fused CA is caused by patch-wise blending of source and target 2D semantics, which introduces ambiguous conditions that mislead the generation model. Specifically, the blended keys and values in cross-attention are derived from patch-wise DINOv2 [52] features. However, spatially aligned patches between source and target images often lack semantic correspondence, leading to distorted outputs. To verify this hypothesis, we visualize the cross-attention maps of head SLAT features (marked by red stars) in Fig. 5. The attention maps from vanilla CA (first two columns) correctly focus on the head regions in the input conditions (indicated by pink stars). In contrast, KV-Fused CA (third column) erroneously attends to background regions (highlighted by an orange box), subsequently using semantically mismatched features to guide head SLAT generation—resulting in local structural distortions. To address this issue, we propose Morphing Cross-Attention (MCA) (Fig. 3-(b)). Instead of blending keys and values prior to the attention computation, MCA computes separate attention outputs for the source and target objects and then combines them to produce the final feature $f^{n}_{MCA}$:
@@ 3.5 Temporal-Fused Self-Attention (TFSA)
Although MCA yields structurally coherent and visually plausible deformations, temporal smoothness remains limited due to the frame-wise independence inherent in standard generative pipelines. To address this limitation—and building on insights from KV-Fused SA discussed in Sec. 3.3—we propose Temporal-Fused Self-Attention (TFSA). As illustrated in Fig. 3-(c), TFSA promotes temporally consistent 3D morphing by incorporating features from previously generated frames into the self-attention mechanism. Specifically, when generating the $n$ frame, TFSA fuses the keys and values of the current frame with those from the immediately preceding frame to produce the output feature $f^{n}_{TFSA}$:
@@ 4.2 Evaluation
Qualitative Results. As shown in Fig. 7, MorphAny3D generates smooth, high-quality 3D morphing sequences across diverse categories, outperforming all baselines. Matching-based methods (e.g., 3DInterp, SLATInterp, MorphFlow) yield smooth, but often implausible or semantically incoherent transitions. 2D-first approaches (DiffMorpher, FreeMorph) leverage strong 2D priors for structural plausibility but suffer from poor 3D temporal consistency, leading to jerky sequences. DirectInterp’s feature fusion is ill-suited to SLAT, resulting in suboptimal deformations. In contrast, our method integrates SLAT features with Trellis’s 3D generative priors to produce temporally smooth, semantically meaningful, and visually faithful morphs. For example, in the elephant-to-excavator case (Fig. 7), it implicitly aligns the trunk with the boom, creating a coherent hybrid that seamlessly blends both concepts.
@@ 5 Conclusion
Limitations and Future Work. MorphAny3D inherits Trellis’s limitations and may exhibit artifacts on extremely fine structures (see red boxes in Fig. 11). Future work could adopt stronger 3D generative backbones [28] or enhance SLAT representations to capture finer geometric details.

### v3

@@ Abstract
Figure 1: MorphAny3D enables smooth, semantically coherent, and aesthetically pleasing transitions from source objects (first column) to target objects (last column), even when they share no semantic or visual relationship (e.g., bee to biplane in the second row).
@@ 1 Introduction
  • We propose MorphAny3D, the first training-free 3D morphing framework based on Structured Latent (SLAT), capable of generating smooth and semantically coherent deformations across diverse 3D object categories.
@@ 3.3 SLAT Fusion Patterns in Attention Mechanism
From these results, we draw three key observations: (1) KV-Fused CA significantly enhances the structural and semantic plausibility of 3D morphing by fusing 2D conditions from the source and target objects within cross-attention (the lowest FID). However, it fails to eliminate localized implausible artifacts, as shown in the orange box of Fig. 4-(b). (2) KV-Fused SA effectively improves the smoothness and continuity of the morphing sequence by aggregating 3D latent features from the source and target objects within self-attention (the lowest PPL). (3) Yet, simultaneously applying KV-fused CA and KV-fused SA—while aiming to maximize both plausibility and smoothness—compromises the plausibility of the resulting deformation (see Fig. 4-(d)). These results indicate that while SLAT feature blending benefits morphing, naive combination harms the plausibility-smoothness trade-off. To address this, we propose two tailored modules—Morphing Cross-Attention and Temporal-Fused Self-Attention—that preserve the advantages of fusion while avoiding its pitfalls.
@@ 3.4 Morphing Cross-Attention (MCA)
We hypothesize that the local irrational structure in KV-Fused CA is caused by patch-wise blending of source and target 2D conditions, which introduces ambiguous semantics that mislead the generation model. Specifically, the blended keys and values in cross-attention are derived from patch-wise DINOv2 [52] features. However, spatially aligned patches between source and target images often lack semantic correspondence, leading to distorted outputs. To verify this hypothesis, we visualize the cross-attention maps of head SLAT features (marked by red stars) in Fig. 5. The attention maps from vanilla CA (second and third columns) correctly focus on the head regions in the input conditions (indicated by pink stars). In contrast, KV-Fused CA (fourth column) erroneously attends to background regions (highlighted by an orange box), subsequently using semantically mismatched features to guide head SLAT generation—resulting in local structural distortions. To address this issue, we propose Morphing Cross-Attention (MCA) (Fig. 3-(b)). Instead of blending keys and values prior to the attention computation, MCA computes separate attention outputs for the source and target objects and then combines them to produce the final feature $f^{n}_{MCA}$:
@@
Figure 5: Attention maps visualization for different attention mechanisms. Red stars denote head SLAT features; pink stars mark their corresponding input regions. Orange boxes highlight KV-Fused CA’s incorrect attention focus. MCA preserves correct, semantically consistent attention and avoids KV-Fused CA’s artifacts shown in Fig. 4-(b).
Figure 6: t-SNE visualization of attention features for different attention mechanisms in the SS Stage (left) and SLAT Stage (right). Obviously, MCA exhibits a stable, smooth feature trajectory, whereas KV-Fused CA’s is erratic and interrupted.
Furthermore, we verified MCA’s efficacy via feature distribution analysis. Fig. 6 shows t-SNE [71] visualizations of average output features from vanilla CA, MCA, and KV-Fused CA across all blocks and diffusion timesteps in the SS and SLAT stages (green/purple deepen during morphing). Notably, MCA exhibits a stable, smooth trajectory, whereas KV-Fused CA yields an erratic, interrupted path. Combined with Fig. 5, this reveals that naive KV fusion disrupts conditions’ semantic consistency, causing unstable 3D morphing (interrupted purple trajectory) and chaotic outputs (purple arrow). In contrast, MCA preserves accurate, semantically consistent attention maps, enabling smooth, plausible 3D deformations (stable green trajectory and arrow).
@@ 3.5 Temporal-Fused Self-Attention (TFSA)
Although MCA yields structurally coherent and visually plausible deformations, temporal smoothness remains limited due to the frame-wise independence inherent in standard 3D generators. To address this limitation—and building on insights from KV-Fused SA discussed in Sec. 3.3—we propose Temporal-Fused Self-Attention (TFSA). As illustrated in Fig. 3-(c), TFSA promotes temporally consistent 3D morphing by incorporating features from previously generated frames into the self-attention mechanism. Specifically, when generating the $n$ frame, TFSA fuses attention outputs from current and previous frames’ keys and values to produce the output $f^{n}_{TFSA}$:
@@ 4.2 Evaluation
Qualitative Results. As shown in Fig. 8, MorphAny3D generates smooth, high-quality 3D morphing sequences across diverse categories, outperforming all baselines. Matching-based methods (e.g., 3DInterp, SLATInterp, MorphFlow) yield smooth, but often implausible or semantically incoherent transitions. 2D-first approaches (DiffMorpher, FreeMorph) leverage strong 2D priors for structural plausibility but suffer from poor 3D temporal consistency, leading to jerky sequences. DirectInterp’s feature fusion is ill-suited to SLAT, resulting in suboptimal deformations. In contrast, our method unleashes the potential of 3D generative priors in SLAT to produce temporally smooth, semantically meaningful, and visually faithful morphs. For example, in the elephant-to-excavator case (Fig. 8), it implicitly aligns the trunk with the boom, creating a coherent hybrid that seamlessly blends both concepts.
@@ 5 Conclusion
Limitations and Future Work. MorphAny3D inherits Trellis’s limitations and may exhibit artifacts on extremely fine structures (see red boxes of case 1 in Fig. 12). Future work could adopt stronger 3D generative backbones [28] or enhance SLAT representations to capture finer geometric details. Additionally, our orientation correction strategy may miss yaw-symmetric objects’ rotations (see case 2 in Fig. 12). Finally, the high runtime restricts its application; we plan to investigate acceleration methods like KV cache.
Acknowledgement. This work was supported by the NSFC under Grant No. 62376121, Basic Research Program of Jiangsu under Grant No. BK20251999, Gusu Innovation Leading Talent Program under Grant No. ZXL2025319, and Jiangsu Provincial Science & Technology Major Project under Grant No. BG2024042.